11) Determinc & when / = 0.50 A and R= 12 A) 12 V B) 24 V D) 13 V C) 15 V
Answer #1

current in 2R = I*2R/(2R+2R)

0.5 = I/2

I = 1A

current in the battery = 1A

2R and 2R are in parallel

R = (2R)*(2R) / (2R + 2R) = R

now R and R are in series

R(equi) = R + R = 2R = 2*12 = 24 ohm

e = IR = 1*24 = 24 V

so option B) is the correct answer